I need help for this system , can you please help me? I will really appreciate it
1/x + 1/y =6
2x+y=3
I've tried to subtract y and then to solve the system but didnt work .then I 've tried to notate 1/x with a and 1/y with b, but I got stack because I didn't know what should I do with the second equation .
thank you for your help.
daniela

Answer by scott8148(6628)

from 2nd equation ___ y = 3-2x

substituting ___ (1/x) + [1/(3-2x)] = 6

3-2x + x = 6(x)(3-2x) ___ 3-2x + x = 18x - 12x^2

0 = 12x^2 - 19x + 3

use quadratic formula to find the x-values , then substitute to find the corresponding y-values